# Tight concentration inequalities for quantum adversarial setups exploiting permutation symmetry

### v2

@@ II.3 Concentration inequality for an independent and identical measurement on an adversarial state
$R=\left\{\bm{p}\in{\cal P}:\frac{\bm{w}}{\|\bm{w}\|}\cdot\bm{p}=-\frac{\bm{u}\cdot\bm{v}}{\sqrt{k}\|\bm{w}\|}\right\}.$ (133)
Since $R^{\prime}$ shifts $R$ by $\sqrt{2}/n$ towards $\bm{p}_{1}$, i.e., along $\bm{w}$, we have
$R^{\prime}$ =$=\left\{\bm{p}\in{\cal P}:\frac{\bm{w}}{\|\bm{w}\|}\cdot\bm{p}=\frac{\sqrt{2}}{n}-\frac{\bm{u}\cdot\bm{v}}{\sqrt{k}\|\bm{w}\|}\right\}$ (134)
=$=\left\{\bm{p}\in{\cal P}:\bm{u}\cdot\bm{p}=\frac{\sqrt{2}\|\bm{w}\|}{n}\right\}.$ (135)
Let ${\cal R}^{\prime}$ be defined as ${\cal R}^{\prime}\coloneqq\{\bm{p}\in{\cal P}:\bm{u}\cdot\bm{p}\leq\sqrt{2}\|\bm{w}\|/n\}$, whose boundary is given by $R^{\prime}$. Then, the set $\tilde{{\cal R}}$ of lattice points mentioned above can be given by ~$\tilde{\cal R}=\tilde{\cal P}\cap\bigl{(}({\cal R}^{\prime}\setminus{\cal R})\cup R\bigr{)}$. Now, we prove that every path in $\tilde{\cal P}$ from $(1,0,\ldots,0)$ to $(0,n_{2},\ldots,n_{k})/n$ (for any possible combination $(n_{2},\ldots,n_{k})$ with $\sum_{j=2}^{k}n_{j}=n$) by repeatedly subtracting $1/n$ from the first entry and adding $1/n$ to one of the other entries hits at least one element in $\tilde{\cal R}$. We define the neighbors of a lattice point $\tilde{\bm{m}}$ in $\tilde{\cal P}$ if they can be reached by subtracting $1/n$ from one of the entries and adding $1/n$ to another entry of $\tilde{\bm{m}}$. Then, the Euclidean distance between neighboring lattice points is $\sqrt{2}/n$. If there exists a path from $(1,0,\ldots,0)$ to $(0,n_{2},\ldots,n_{k})/n$ that goes across the region sandwiched by $R$ and $R^{\prime}$ (i.e., $({\cal R}^{\prime}\setminus{\cal R})\cup R$) without hitting any element in $\tilde{\cal R}$, then that means there exists a pair of lattice points that are neighbors but in opposite sides of the region. But this contradicts the fact that two parallel hyperplanes $R$ and $R^{\prime}$ are separated by $\sqrt{2}/n$. This completes the proof by contradiction.
Now, let us consider the following convex mixture $\sigma$ of i.i.d. probability distribution over ${\cal X}^{\times n}$:
$\sigma(\bm{j})$ $=\sum_{\begin{subarray}{c}\bm{m}=(n_{1},\ldots,n_{k})\in Z_{n}^{k}\\ :\bm{m}/n\in\tilde{\cal R}\end{subarray}}\mathrm{P}_{\tilde{\rho}}(\bm{m})\prod_{i=1}^{n}\frac{n_{j_{i}}}{n}$ (136)
@@
which is not necessarily normalized. Let $\mathrm{Q}_{\sigma}:Z_{n}^{k}\to[0,1]$ be a probability mass function associated with $\sigma$ defined as
$\mathrm{Q}_{\sigma}(\bm{m})$ $\coloneqq\sum_{\bm{j}\in{\cal T}(\bm{m})}\sigma(\bm{j})$ (138)
@@
From Eqs. (12) and (83), we have, for any $\bm{m}$ with $\bm{m}/n\in\tilde{\cal R}$,
@@
In the following, we inductively prove that $\mathrm{P}_{\tilde{\rho}}(\bm{m})\leq f_{c}(n,k)\mathrm{Q}_{\sigma}(\bm{m})$ holds for any $\bm{m}$ with $\bm{m}/n\in({\cal P}\setminus{\cal R})\cup R$. Let us define ~$\tilde{\cal R}^{(0)}\coloneqq\tilde{\cal R}$. Then, for any frequency distribution $\bm{m}$ with its type in ~$\tilde{\cal R}^{(0)}$, Eq. (140) holds. Let $\bm{l}^{\prime}$ be an arbitrary frequency distribution such that $\bm{l}^{\prime}[1]/n\in\tilde{\cal P}\setminus\tilde{\cal R}^{(0)}$ and $\bm{l}^{\prime}[j]/n\in\tilde{\cal R}^{(0)}$ for $j=2,\ldots,k$. (Note that $\bm{l}^{\prime}[1]/n\in{\cal P}\setminus{\cal R}^{\prime}$ automatically holds if the above conditions are satisfied.) Such a sequence $\bm{l}^{\prime}$ always exists; if there does not exist such a sequence, then that means there exists $j$ such that $\bm{l}^{\prime}[1]/n\in\tilde{\cal P}\setminus\tilde{\cal R}^{(0)}$ and $\bm{l}^{\prime}[j]/n\notin\tilde{\cal R}^{(0)}$. If $\bm{l}^{\prime}[j]/n\in(\tilde{\cal P}\setminus\tilde{\cal R}^{(0)})\cap({\cal P}/{\cal R})$, then we can find a sequence $\bm{l}^{\prime\prime}$ that satisfies the imposed condition by reducing the first entry of $\bm{l}^{\prime}$ and add it to the $j$-th entry. The other possibility is that $\bm{l}^{\prime}[j]/n\in{\cal R}\setminus R$, but this implies that we can find a path that goes from $\bm{p}_{1}$ through $\bm{l}^{\prime}[1]/n\in\tilde{\cal P}\setminus\tilde{\cal R}^{(0)}$ and $\bm{l}^{\prime}[j]/n\in{\cal R}\setminus R$ to $(0,n_{2},\ldots,n_{k})/n$ (with $\sum_{j=2}^{k}n_{j}=n$) without intersecting ~$\tilde{\cal R}^{(0)}$, which contradicts the fact that every path in $\tilde{\cal P}$ from $\bm{p}_{1}$ to $(0,n_{2},\ldots,n_{k})/n$ hits at least one element in $\tilde{\cal R}$. From the constraint on $\mathrm{P}_{\tilde{\rho}}(\cdot)$ in Eq. (130), we have
$\frac{\mathrm{P}_{\tilde{\rho}}(\bm{l}^{\prime}[1])}{|{\cal T}(\bm{l}^{\prime}[1])|}-\biggl{(}\sum_{j=2}^{k}b_{j}\frac{\mathrm{P}_{\tilde{\rho}}(\bm{l}^{\prime}[j])}{|{\cal T}(\bm{l}^{\prime}[j])|}\biggr{)}\leq 0.$ (141)
@@
Furthermore, since $\sigma$ is a mixture of the i.i.d. probability distributions and is thus invariant under the permutation of $n$ outcomes, it has the same form as the right-hand side of Eq. (125) with $\mathrm{P}_{\tilde{\rho}}$ replaced with $\mathrm{Q}_{\sigma}$. Considering again a conditional probability mass function of the $n$-th outcome $i_{n}$ conditioned on any sequence $\bm{i}^{\prime}=(i_{1},\ldots,i_{n-1})$ of the other $n-1$ outcomes of a particular sequence $\bm{i}=(i_{1},\ldots,i_{n})$ for $\sigma$, we notice that the resulting conditional probability mass function is an $\bm{i}^{\prime}$ dependent mixture of the probability distributions in $\tilde{\cal R}$ since $\sigma$ is a mixture of the i.i.d. probability distributions in $\tilde{\cal R}$. Combining these facts, we have
@@
$\begin{split}&\left(\sum_{j=1}^{k}\frac{\mathrm{Q}_{\sigma}(\bm{l}^{\prime}[j])}{|{\cal T}(\bm{l}^{\prime}[j])|}\right)^{-1}\left(\frac{\mathrm{Q}_{\sigma}(\bm{l}^{\prime}[1])}{|{\cal T}(\bm{l}^{\prime}[1])|},\ldots,\frac{\mathrm{Q}_{\sigma}(\bm{l}^{\prime}[k])}{|{\cal T}(\bm{l}^{\prime}[k])|}\right)\\ &\qquad\qquad\qquad\in({\cal R}^{\prime}\setminus{\cal R})\cup R\subset({\cal P}\setminus{\cal R})\cup R,\end{split}$ (142)
and thus have
$\frac{\mathrm{Q}_{\sigma}(\bm{l}^{\prime}[1])}{|{\cal T}(\bm{l}^{\prime}[1])|}-\biggl{(}\sum_{j=2}^{k}b_{j}\frac{\mathrm{Q}_{\sigma}(\bm{l}^{\prime}[j])}{|{\cal T}(\bm{l}^{\prime}[j])|}\biggr{)}\geq 0.$ (143)
@@
$\mathrm{P}_{\tilde{\rho}}(\bm{l}^{\prime}[1])\leq f_{c}(n,k)\mathrm{Q}_{\sigma}(\bm{l}^{\prime}[1]).$ (144)
Thus, we have the same inequality as Eq. (140) for $\bm{l}^{\prime}[1]$. We repeatedly apply the above argument by replacing ~$\tilde{\cal R}^{(0)}$ with ~$\tilde{\cal R}^{(1)}\coloneqq\tilde{\cal R}^{(0)}\cup\{\bm{l}^{\prime}[1]/n\}$ and obtains $\mathrm{P}_{\tilde{\rho}}(\bm{m})\leq f_{c}(n,k)\mathrm{Q}_{\sigma}(\bm{m})$ for any frequency distribution $\bm{m}$ with $\bm{m}/n\in({\cal P}\setminus{\cal R})\cup R$. We also have $\mathrm{P}_{\tilde{\rho}}(\bm{m})\leq f_{c}(n,k)\,\mathrm{Mult}_{\bm{m}/n}(\bm{m})$ for any $\bm{m}/n\in{\cal R}$ from Eq. (12) and (83). Thus, for any $\bm{m}\in Z_{n}^{k}$, we have
$\mathrm{P}_{\tilde{\rho}}(\bm{m})\leq\sum_{\bm{m}^{\prime}:\bm{m}^{\prime}/n\in{\cal R}^{\prime}}f_{c}(n,k)\mathrm{P}_{\tilde{\rho}}(\bm{m}^{\prime})\,\mathrm{Mult}_{\bm{m}^{\prime}/n}(\bm{m}).$ (145)
This inequality immediately implies that for any subset $A$ of $Z_{n}^{k}$, we have
$\sum_{\bm{m}\in A}\mathrm{P}_{\tilde{\rho}}(\bm{m})\leq\max_{\bm{q}\in{\cal R}^{\prime}}\sum_{\bm{m}\in A}f_{c}(n,k)\,\mathrm{Mult}_{\bm{q}}(\bm{m}).$ (146)
@@
$D(\bm{p}\|\bm{q})\coloneqq\begin{cases}\mathrm{Tr}[\bm{p}\ln\bm{p}-\bm{p}\ln\bm{q}]&\text{supp}(\bm{p})\subseteq\text{supp}(\bm{q})\\ +\infty&\text{Otherwise}\end{cases},$ (147)
where $\text{supp}(\bm{r})$ denotes the support of the (discrete) probability distribution $\bm{r}$. Let ${\cal A}$ be a closed convex subset of ${\cal P}$. Then, the probability that the type $\bm{m}/n$ of the outcomes of the independent and identical measurements lie in ${\cal A}$ is bounded from above by
@@
$\leq\max_{\bm{p}\in{\cal A}}\max_{\bm{q}\in{\cal R}^{\prime}}f_{c}(n,k)\exp[-nD(\bm{p}\|\bm{q})],$ (149)
where we used the refined version of Sanov’s inequality [2, 3] in the second inequality. Combining this with Eqs. (123) and (124), we prove Theorem 1.∎
@@
III Numerical comparison of the performance with other concentration inequalities
The obtained bound in the previous section has only a polynomial factor compared to the i.i.d. case, and thus it is asymptotically tight. However, what one may have an interest in a practical setup is tightness in the finite-size regime. This is especially important for applications such as QKD. Here, we compare its tightness with other known concentration inequalities that can deal with non-commutative observables. One famous example is Azuma’s inequality [26], which is conventionally used in the so-called phase error correction approach to prove the security of QKD protocols. Azuma’s inequality does not require permutation invariance unlike Theorem 1 and can be applied to arbitrary correlated random variables in combination with the so-called Doob decomposition theorem [30]. Various generalizations of Azuma’s inequality have been studied [10, 11], most of which have tightened the inequality by using the bound on the higher moments of the true distribution. Recently, Kato has succeeded in tightening Azuma’s inequality by using unconfirmed knowledge, i.e., a priori knowledge on the expectation value of random variables [12]. The inequality holds even when a priori knowledge is false, and it tightens the inequality to the level of i.i.d. random sampling when a priori knowledge is true, which is extremely useful in applications such as QKD. Since our method can also be applied to adversarial setups such as QKD, we compare the performance of our new method with these two conventionally used bounds.

### v3

@@ 2.3 Concentration inequality for an independent and identical measurement on an adversarial state
We prove that every path in $\tilde{\cal P}$ from $\bm{p}_{1}=(1,0,\ldots,0)$ to $(0,n_{2},\ldots,n_{k})/n\in{\cal R}$ (for any possible combination $(n_{2},\ldots,n_{k})$ with $\sum_{j=2}^{k}n_{j}=n$) by repeatedly subtracting $1/n$ from the first entry and adding $1/n$ to one of the other entries hits at least one element in $\tilde{\cal R}$. We define the neighbors of a type $\bm{p}\in\tilde{\cal P}$ to be those types that can be reached by subtracting $1/n$ from one entry of $\bm{p}$ and adding $1/n$ to another entry of $\bm{p}$. Then, the Euclidean distance between neighboring types is $\sqrt{2}/n$. If there exists a path from $(1,0,\ldots,0)$ to $(0,n_{2},\ldots,n_{k})/n$ that does not hit any element of $\tilde{\cal R}$, then that means there exists a pair $(\bm{p},\bm{p}^{\prime})$ of neighboring probability vectors such that $\bm{p}\in{\cal P}\setminus{\cal R}^{\prime}$ and $\bm{p}^{\prime}\in{\cal R}$. But this contradicts the fact that any $\bm{q}\in{\cal P}$ with $\|\bm{q}-\bm{p}^{\prime}\|\leq\sqrt{2}/n$ is contained in ${\cal R}^{\prime}$. This completes the proof by contradiction.
Now, let us consider the following subnormalized convex mixture $\sigma$ of i.i.d. probability distributions over ${\cal X}^{\times n}$:
$\sigma(\bm{i})$ $=\sum_{\begin{subarray}{c}\bm{m}=(n_{1},\ldots,n_{k})\in Z_{n}^{k}\\ :\bm{m}/n\in\tilde{\cal R}\end{subarray}}\mathrm{P}_{\tilde{\rho}}(\bm{m})\prod_{l=1}^{n}\frac{n_{i_{l}}}{n}$ (139)
@@
which is subnormalized since $\sum_{\bm{m}\in Z_{n}^{k}:\bm{m}/n\in\tilde{\cal R}}\mathrm{P}_{\tilde{\rho}}(\bm{m})$ may not be unity in general. Let $\mathrm{Q}_{\sigma}:Z_{n}^{k}\to[0,1]$ be a probability mass function associated with $\sigma$ defined as
$\mathrm{Q}_{\sigma}(\bm{m})$ $\coloneqq\sum_{\bm{i}\in{\cal T}(\bm{m})}\sigma(\bm{i})$ (141)
@@
From Eqs. (14) and (85), we have, for any $\bm{m}\in Z_{n}^{k}$ with $\bm{m}/n\in\tilde{\cal R}$,
@@
In the following, we inductively prove that $\mathrm{P}_{\tilde{\rho}}(\bm{m})\leq f_{c}(n,k)\mathrm{Q}_{\sigma}(\bm{m})$ holds for any $\bm{m}\in Z_{n}^{k}$ with $\bm{m}/n\in{\cal P}\setminus{\cal R}$. Let us define ~$\tilde{\cal R}^{(0)}\coloneqq\tilde{\cal R}$. Then, for any frequency distribution $\bm{m}$ with its type in ~$\tilde{\cal R}^{(0)}$, Eq. (143) holds. Let $\bm{l}\in Z^{k}_{n-1}$ be an arbitrary frequency distribution such that $\bm{l}[1]/n\in\tilde{\cal P}\setminus\tilde{\cal R}^{(0)}$ and $\bm{l}[j]/n\in\tilde{\cal R}^{(0)}$ for $j=2,\ldots,k$. (Note that $\bm{l}[1]/n\in{\cal P}\setminus{\cal R}^{\prime}$ automatically holds if the above conditions are satisfied.) Such a sequence $\bm{l}$ always exists; if there does not exist such a sequence, then that means there exists $j$ such that $\bm{l}[1]/n\in\tilde{\cal P}\setminus\tilde{\cal R}^{(0)}$ and $\bm{l}[j]/n\notin\tilde{\cal R}^{(0)}$. If $\bm{l}[j]/n\in\tilde{\cal P}\setminus\tilde{\cal R}^{(0)}$, then we can find a sequence $\bm{l}^{\prime}$ that satisfies the imposed condition by reducing the first entry of $\bm{l}$ by a positive integer multiple of $1/n$ and add it to the $j$-th entry. The other possibility is that $\bm{l}[j]/n\in{\cal R}$, but this implies that we can find a path that goes from $\bm{p}_{1}$ through $\bm{l}[1]/n\in\tilde{\cal P}\setminus\tilde{\cal R}^{(0)}$ and $\bm{l}[j]/n\in{\cal R}$ to $(0,n_{2},\ldots,n_{k})/n$ (with $\sum_{j=2}^{k}n_{j}=n$) without intersecting ~$\tilde{\cal R}^{(0)}$, which contradicts the fact that every path in $\tilde{\cal P}$ from $\bm{p}_{1}$ to $(0,n_{2},\ldots,n_{k})/n$ hits at least one element in $\tilde{\cal R}$. From the constraint on $\mathrm{P}_{\tilde{\rho}}(\cdot)$ in Eq. (132) and the definition of ${\cal R}$ in Eq. (57), we have
$\frac{\mathrm{P}_{\tilde{\rho}}(\bm{l}[1])}{|{\cal T}(\bm{l}[1])|}-\biggl{(}\sum_{j=2}^{k}b_{j}\frac{\mathrm{P}_{\tilde{\rho}}(\bm{l}[j])}{|{\cal T}(\bm{l}[j])|}\biggr{)}\leq 0.$ (144)
@@
Furthermore, since $\sigma$ is a mixture of the i.i.d. probability distributions and is thus invariant under the permutation of $n$ outcomes, it has the same form as the right-hand side of Eq. (127) with $\mathrm{P}_{\tilde{\rho}}$ replaced with $\mathrm{Q}_{\sigma}$. Considering again a conditional probability mass function of the $n$-th outcome $i_{n}$ conditioned on a sequence $\bm{i}^{\prime}=(i_{1},\ldots,i_{n-1})$ of the other $n-1$ outcomes of a particular sequence $\bm{i}=(i_{1},\ldots,i_{n})$ for $\sigma$, we notice that the resulting conditional probability mass function is an $\bm{i}^{\prime}$ (or in fact $\bm{l}$) dependent mixture of probability distributions in $\tilde{\cal R}$ since $\sigma$ is a mixture of i.i.d. probability distributions in $\tilde{\cal R}$. Combining these facts, we have
@@
$\begin{split}&\left(\sum_{j=1}^{k}\frac{\mathrm{Q}_{\sigma}(\bm{l}[j])}{|{\cal T}(\bm{l}[j])|}\right)^{-1}\left(\frac{\mathrm{Q}_{\sigma}(\bm{l}[1])}{|{\cal T}(\bm{l}[1])|},\ldots,\frac{\mathrm{Q}_{\sigma}(\bm{l}[k])}{|{\cal T}(\bm{l}[k])|}\right)\\ &\qquad\in{\cal R}^{\prime}\setminus{\cal R}\subseteq{\cal P}\setminus{\cal R},\end{split}$ (145)
and thus have from Eq. (131) that
$\frac{\mathrm{Q}_{\sigma}(\bm{l}[1])}{|{\cal T}(\bm{l}[1])|}-\biggl{(}\sum_{j=2}^{k}b_{j}\frac{\mathrm{Q}_{\sigma}(\bm{l}[j])}{|{\cal T}(\bm{l}[j])|}\biggr{)}\geq 0.$ (146)
@@
$\mathrm{P}_{\tilde{\rho}}(\bm{l}[1])\leq f_{c}(n,k)\mathrm{Q}_{\sigma}(\bm{l}[1]).$ (147)
Thus, we have the same inequality as Eq. (143) for $\bm{l}[1]$. We repeatedly apply the above argument by replacing ~$\tilde{\cal R}^{(0)}$ with ~$\tilde{\cal R}^{(1)}\coloneqq\tilde{\cal R}^{(0)}\cup\{\bm{l}[1]/n\}$ and so on, and finally obtains $\mathrm{P}_{\tilde{\rho}}(\bm{m})\leq f_{c}(n,k)\mathrm{Q}_{\sigma}(\bm{m})$ for any frequency distribution $\bm{m}\in Z_{n}^{k}$ with $\bm{m}/n\in{\cal P}\setminus{\cal R}$ by induction.
From Eq. (14) and (85), we also have
$\mathrm{P}_{\tilde{\rho}}(\bm{m})\leq\sum_{\bm{m}^{\prime}:\bm{m}^{\prime}/n\in{\cal R}}f_{c}(n,k)\,\mathrm{Mult}_{\bm{m}^{\prime}/n}(\bm{m})$ (148)
for any $\bm{m}/n\in{\cal R}$. Thus, for any $\bm{m}\in Z_{n}^{k}$, we have
$\begin{split}\mathrm{P}_{\tilde{\rho}}(\bm{m})&\leq\sum_{\bm{m}^{\prime}:\bm{m}^{\prime}/n\in\tilde{\cal R}}f_{c}(n,k)\mathrm{P}_{\tilde{\rho}}(\bm{m}^{\prime})\,\mathrm{Mult}_{\bm{m}^{\prime}/n}(\bm{m})\\ &\quad+\sum_{\bm{m}^{\prime}:\bm{m}^{\prime}/n\in{\cal R}}f_{c}(n,k)\mathrm{P}_{\tilde{\rho}}(\bm{m}^{\prime})\,\mathrm{Mult}_{\bm{m}^{\prime}/n}(\bm{m})\end{split}$ (149)
$=\sum_{\bm{m}^{\prime}:\bm{m}^{\prime}/n\in{\cal R}^{\prime}}f_{c}(n,k)\mathrm{P}_{\tilde{\rho}}(\bm{m}^{\prime})\,\mathrm{Mult}_{\bm{m}^{\prime}/n}(\bm{m}).$ (150)
This inequality immediately implies that for any subset $B$ of $Z_{n}^{k}$, we have
$\sum_{\bm{m}\in B}\mathrm{P}_{\tilde{\rho}}(\bm{m})\leq\max_{\bm{q}\in{\cal R}^{\prime}}\sum_{\bm{m}\in B}f_{c}(n,k)\,\mathrm{Mult}_{\bm{q}}(\bm{m}).$ (151)
@@
$\begin{split}&D(\bm{p}\|\bm{q})\\ &\coloneqq\begin{cases}\mathrm{Tr}[\bm{p}\ln\bm{p}-\bm{p}\ln\bm{q}]&\text{supp}(\bm{p})\subseteq\text{supp}(\bm{q})\\ +\infty&\text{Otherwise}\end{cases},\end{split}$ (152)
where $\text{supp}(\bm{r})$ denotes the support of the (discrete) probability distribution $\bm{r}$. Then, the following theorem is known [3].
Theorem 4 (Refined Sanov’s theorem for a closed convex set [3]).
Let ${\cal A}$ be a non-empty closed convex subset of the $(k-1)$-simplex ${\cal P}$ of probability distributions. Let $\bm{q}\in{\cal P}$ be a probability distribution, and let $Z_{n}^{k}$ be as defined in Eq. (5). Then, for a multinomial distribution $\mathrm{Mult}_{\bm{q}}$ of $\bm{q}$, one has
$\sum_{\bm{m}\in Z_{n}^{k}:\bm{m}/n\in{\cal A}}\mathrm{Mult}_{\bm{q}}(\bm{m})\leq\max_{\bm{p}\in{\cal A}}\exp[-nD(\bm{p}\|\bm{q})].$ (153)
(In the above theorem, if ${\cal A}$ is not restricted to a closed convex subset, then a polynomial factor is multiplied to the right-hand side, which may be more conventional [2, 32]. However, for our purpose of applying it to quantum measurement scenario, we typically consider a subset with an affine boundary obtained through an operator inequality, and thus the above mentioned restriction on ${\cal A}$ may not lose generality.) Combining the above theorem with Eq. (151), we obtain
@@
$\leq\max_{\bm{p}\in{\cal A}}\max_{\bm{q}\in{\cal R}^{\prime}}f_{c}(n,k)\exp[-nD(\bm{p}\|\bm{q})].$ (155)
Note that the simultaneous maximization over $\bm{p}\in{\cal A}$ and $\bm{q}\in{\cal R}^{\prime}$ is a convex optimization problem, and thus the order of maximization does not matter. Combining this with Eqs. (125) and (126), we prove Theorem 3.∎
@@
3 Numerical comparison of the performance with other concentration inequalities
The obtained bound in the previous section has only a polynomial factor compared to the i.i.d. case, and thus it is asymptotically tight. However, what one may have an interest in a practical setup is tightness in the finite-size regime. This is especially important for applications such as QKD. Here, we compare its tightness with other known concentration inequalities that can deal with non-commutative observables. One famous example is Azuma’s inequality [28], which is conventionally used in the so-called phase error correction approach to prove the security of QKD protocols. Azuma’s inequality does not require permutation invariance unlike Theorem 1 and can be applied to arbitrary correlated random variables in combination with the so-called Doob decomposition theorem [33]. Various generalizations of Azuma’s inequality have been studied [12, 13], most of which have tightened the inequality by using the bound on the higher moments of the true distribution. Recently, Kato has succeeded in tightening Azuma’s inequality by using unconfirmed knowledge, i.e., a priori knowledge on the expectation value of random variables [14]. The inequality holds even when a priori knowledge is false, and it tightens the inequality to the level of i.i.d. case for a binary random variable when a priori knowledge is true, which is extremely useful in applications such as QKD. Since our method can also be applied to adversarial setups such as QKD, we compare the performance of our new method with these two conventionally used bounds.